OpenStudy (anonymous):

Which of the following is a factor of 2x*4 + 22x*3 + 60x*2? a. 2x*3 b. x*4 c. x+4 d. x+5

OpenStudy (anonymous):

@pooja195

OpenStudy (danjs):

First, what is the biggest factor of the coefficients, 2 , 22 , 60 ?

OpenStudy (anonymous):

60

OpenStudy (danjs):

no, you can not take a multiple of 60 out of 2 or 22

OpenStudy (danjs):

2 is the largest thing that can divide evenly into all three of those

OpenStudy (danjs):

so you can say \[2x^4+22x^3+60x^2 = 2*(x^4+11x^3+30x^2)\]

OpenStudy (anonymous):

yeah

OpenStudy (danjs):

now, since you have just x variables, the greatest one you can pull out of all three will be the smallest power, only if every term has x, like here

OpenStudy (danjs):

\[2*(x^4+11x^3+30x^2) = 2x^2*(x^2+11x+30)\]

OpenStudy (danjs):

see that?

OpenStudy (danjs):

Really are just multiplying the thing by x^2 / x^2 or 1, and keeping the numerator out front, and dividing everything by the x^2

OpenStudy (danjs):

Now since 2x^2 is not an answer, look further, factor the remaining ( ) part